Intro Samsung has sales of $45,000 and a net income of $20,000. Total assets are $46,000 and total equity is $20,000. The consumer electronics industry as a whole has the following characteristics: Ratio Profit margin Total asset turnover 0.941 Equity multiplier 1.8 Return on equity 0.232 Value 0.137 B Attempt 1/10 for 8 pts. Part 1 What is Samsungs profit margin? 3+ decimals Submit 11 E] -Attempt 1/10 for 8 pts. Part 2 What is Samsungs total asset turnover? 2+ decimals
media%2F9f1%2F9f1abf03-6eed-4194-83bd-23
Answer #1

1) profit margin = net income / sale

= 20000 / 45000

= 0.444

2) total asset turnover = sales / total assets

= 45000 / 46000

= 0.978

3) Equity multiplier = total assets / Total equity

= 46000 / 20000

= 2.3

4) Return on equity = net income / Total equity

= 20000 / 20000

= 1

5) higher profit margin , the higher net income,the greater will be ROE. Greater ROE means company is reinvesting its profit into its buisness to generate higher return